solve each equation.
2/y+ 5/2 = 4 - 2/3y


Answer by totxone(2)

2/y + 2/3y = 4 - 5/2 >>> Bring the like terms together
1/y(2+2/3) = 4 - 2 1/2 >>> took 1/y as common and wrote 5/2 as mixed fraction
1/y(2 2/3) = 1 1/2 >>> added the fractions on both sides
1/y(8/3) = 3/2 >>> converted mixed number to improper fraction
8/3y = 3/2
2X8 = 3 X 3y >>> cross multiply
9y = 16 >>> multiplied the constants

y = 16/9 or 1 7/9